OpenStudy (anonymous):

If a man earns $420 interest in one year at 7%, what was the amount of money invested? $29.40 $60.00 $600.00 $6,000.00

OpenStudy (anonymous):

B?

OpenStudy (akashdeepdeb):

Interest = principal * rate * time/100 Thus 420 * 100 = principal * 7 * 1 [SHIFT THE 100 AND WE KNOW TIME = 1YEAR] Thus principal = $6000 Thus the option is D) $6000

OpenStudy (anonymous):

thanks i thought it couldnt have been that small of a number

OpenStudy (akashdeepdeb):

:) Kudos?

OpenStudy (anonymous):

Yep, again kudos :-)

OpenStudy (uri):

Interest=capital*rate 420=capital*0.7 capital=420/0.7=600 C